A photon of wavelength "l" is scattered from an electron which is at rest. The wavelength shift is "3l". The angle of scattering is 60degrees and the angle at which the electron recoiled is "b". The value of tan(b) is ?
(electron speed is much smaller than the speed of light)

Grade:12

1 Answers

Aziz Alam IIT Roorkee
9 years ago

PLease llok into the solution
We have:
P_{x_{i}}=P_{x_{f}}
\frac{h}{\lambda}=\frac{h}{8\lambda}+P_e

P_e=\frac{7h}{8}\frac{1}{\lambda} Momentum along x-axis of electron

P_e=\frac{\sqrt3}8 \frac{\lambda}{h}] Momentum along y-axis of electron

hence the value of tan
tan(b)=\frac{\sqrt3}7